# Language Model Preference Evaluation with Multiple Weak Evaluators

### v3

@@ 7 Experiments on Instruct Tuning
From Table 3, we observe that GED consistently outperforms all baseline methods, demonstrating its effectiveness in selecting high-quality responses when multiple answers are available for a given instruction. When faced with multiple responses $y_{1},y_{2},\ldots,y_{n}$ for a given instruction $x$, the Random selection method can have a detrimental impact, especially when the quality of the responses is inconsistent. This effect is most evident with the Mistral-7B, where Random selection actually performs worse than the Origin, indicating that randomly chosen data points can introduce noise and degrade the model’s performance. Moreover, we find that simply selecting the longest response does not always lead to the best outcomes. While longer responses may provide more detailed answers, they are not necessarily better in terms of quality, particularly when both high-quality and low-quality answers exist for the same question. This is reflected in the results where the Longest method underperforms compared to both ContraSolver and GED, emphasizing that response length alone is not always a reliable criterion. From Figure 4, we observe similar trends as in Table 3. GED consistently outperforms all baselines across various datasets, demonstrating its effectiveness in selecting high-quality responses. Notably, in AlpacaEval and Self-Instruct, the Random baseline performs worse than the Origin model, highlighting that when response quality varies significantly, poor selection can degrade model performance. In contrast, GED leverages preference graphs and denoising techniques to filter out low-quality responses, ensuring more robust and reliable performance, particularly in settings with inconsistent responses, as it removes evaluation noise and leads to more robust performance.
@@
Table 4: Performance comparison on nuanced quality metrics (%). GED outperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful. and Validity metrics.
Method Factuality Relevance Coherence Inform. Helpful. Validity Avg.
@@
GED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
8 Evaluating GED on More Metrics
@@
To assess GED on more metrics, we evaluated its impact on nuanced quality aspects of LLM outputs, including factuality, relevance, coherence, informativeness, helpfulness, and validity. Since LLM tasks often require subtle judgments, GED’s adaptability is crucial. Following (Chen et al., 2023a), we used Llama3-70B to generate ten candidate responses per query and applied GED for response selection, with Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators. Table 4 shows that GED consistently outperformed individual evaluators and random selection across all metrics. It improved factuality by  5 % points over the best evaluator, while also enhancing relevance, coherence, informativeness, and helpfulness. These results highlight GED’s ability to aggregate preferences effectively, capturing subtle qualities of generated content and improving overall reliability and utility in LLM assessments.
9 Conclusion
In this paper, we presented GED, a framework designed to address inconsistencies in pairwise preference evaluations by LLMs. By employing graph ensemble techniques and denoising, GED reduces cyclic patterns and enhances the reliability of evaluation outcomes. Our theoretical analysis shows that GED can recover the ground truth DAG under reasonable conditions, improving consistency in preference rankings. Extensive experiments across response selection, model ranking, and instruct tuning demonstrate the efficacy of our method. GED consistently outperformed baseline methods in both single-evaluator and multi-evaluator settings, particularly in scenarios where combining small evaluators led to superior results over larger individual evaluators. Future work will explore extending GED to broader evaluation frameworks and applying its principles to more complex decision-making tasks, including multi-agent systems and human-AI interaction.
Impact Statements
This work enhances the reliability of model-based preference evaluation, reducing inconsistencies in AI judgments. It improves fairness and transparency in AI evaluation, benefiting NLP applications and ethical decision-making frameworks.
Appendix A Implementation Details
A.1 Experimental Setup
All experimental procedures were conducted on a machine equipped with an AMD EPYC 7543 32-Core Processor, 512GB memory, 128 CPUs, and four 80GB NVIDIA A800 GPUs. The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Mistral-8×7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-7B-Instruct-v0.3, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct. We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al., 2024; Yu et al., 2023). Each experiment was repeated three times, and the average performance was reported as the final result. Our training script was adapted from the example provided in LlamaFactory (Zheng et al., 2024b)  [^1] https://github.com/hiyouga/LLaMA-Factory/blob/main/examples/train_full/llama3_full_sft_ds3.yaml. The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and he model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler. For generating diverse responses from LLMs, we followed the configuration in Yuan et al. (2024), setting $T=0.7$ and $p=0.9$. For tasks such as AlpacaEval (Dubois et al., 2024b), we used GPT-4-o unless stated otherwise.
A.2 Details of Evaluator Selection Across Different Tasks
Response Selection.
In the response selection task, we evaluated both single models and ensembles of evaluators using the GED method. For single model evaluation, we assessed the standalone performance of Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B on benchmarks such as HumanEval, AlpacaEval, MATH, GSM8k, and GAIA. This served two purposes: first, to establish the baseline performance of smaller models (Llama3-8B, Mistral-7B, Qwen2-7B) that are later combined using GED, and second, to compare against Qwen2-72B, where we tested a baseline approach by randomly selecting one response from the 10 it generated for each question. For single evaluator evaluation, each model (Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B) was also used as an evaluator to rank responses generated by Qwen2-72B. Notably, Qwen2-72B acted as a self-evaluator, selecting the best response from its own generated outputs. Finally, for multiple evaluator evaluation, our GED method combined the evaluations of Llama3-8B, Mistral-7B, and Qwen2-7B to rank responses.
Model Ranking.
For the model ranking task, we selected larger models as evaluators to ensure alignment with rankings produced by GPT-4. Specifically, we used Llama3-70B, Qwen2-72B, Qwen1.5-72B, and Mistral-8×7B as single evaluators. This choice was guided by two factors: performance considerations and practical feasibility. Larger models generally produce more reliable rankings, closely aligning with GPT-4, and the AlpacaEval benchmark, containing 805 tasks, makes the computational cost of using larger models acceptable. In the multiple evaluator setting, our GED method aggregated the evaluations from these four larger models to produce robust and consistent rankings. The combination of these high-capacity models ensures that our approach yields rankings that are both accurate and consistent across tasks.
Instruction Tuning.
In the instruction tuning task, the objective was to perform data selection for model training. For this task, we employed Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators. These models were selected because they balance computational efficiency and evaluation performance, making them suitable for iterative instruction tuning processes. The evaluators’ pairwise preferences were aggregated using GED to identify the most appropriate responses for training. This approach ensured that the selected data pairs reflected a consensus among the evaluators while keeping computational costs manageable.
A.3 Definition of Weak Evaluators
In this work, ”weak evaluators” are defined by their tendency to produce noisy or inconsistent pairwise preferences, not by their overall model capacity. Even advanced models like GPT-4-o can generate preference graphs with cycles, indicating that preference inconsistency is a universal challenge, regardless of model scale. Evaluators such as Llama3-8B, Mistral-7B, Qwen2-7B, and even GPT-4-o are considered weak if their pairwise evaluations exhibit significant noise or conflicts.
A.4 Evaluation Settings: Single Model vs. Single Evaluator
The ”single model” setting evaluates each model’s (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) outputs directly on benchmarks like HumanEval, AlpacaEval, MATH, GSM8k, and GAIA, without selection or modification. In contrast, the ”single evaluator” setting uses these models to select the best response from ten candidates generated by Qwen2-72B, assessing their evaluation capability. The key difference is that the single model setting focuses on generation quality, while the single evaluator setting assesses evaluation ability.
A.5 Definition of Cycle Rate
A.6 Models Used for Ranking
We evaluate model ranking using 30 widely used models from the AlpacaEval dataset, covering diverse architectures and capabilities. Our model set $\mathcal{M}$ includes families such as Llama (Touvron et al., 2023), Vicuna (Zheng et al., 2023), GPT (Achiam et al., 2023), Claude (Anthropic, 2024), Qwen (Yang et al., 2024a), Mistral (Jiang et al., 2023), Yi (Young et al., 2024), and WizardLM (Xu et al., 2023), ensuring a broad assessment. Proprietary models include OpenAI’s GPT series (e.g., gpt-3.5-turbo-0301, gpt-4o-2024-05-13) and Anthropic’s Claude models (e.g., claude-2, claude-3-opus-20240229), serving as strong baselines. Open-source models include multiple versions of Llama (e.g., Meta-Llama-3-8B-Instruct, llama-2-70b-chat-hf) and Qwen (e.g., Qwen1.5-72B-Chat, Qwen2-72B). We also incorporate Mistral models (e.g., Mistral-7B-Instruct-v0.2, Mixtral-8x22B-Instruct-v0.1), reflecting advances in mixture-of-experts architectures. Additional selections include gemini-pro, tulu-2-dpo-70b, oasst-sft-llama-33b, dbrx-instruct, wizardlm-13b, and Yi-34B-Chat, contributing to ranking diversity. This diverse model set enables a comprehensive evaluation of ranking accuracy using AlpacaEval’s ground-truth rankings.
A.7 Dataset
  • UltraFeedback (Cui et al., 2023): UltraFeedback is a large-scale, fine-grained, diverse preference dataset, used for training powerful reward models and critic models. We collect about 64k prompts from diverse resources (including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN). We then use these prompts to query multiple LLMs (see Table for model lists) and generate 4 different responses for each prompt, resulting in a total of 256k samples.
  • HH-RLHF (Bai et al., 2022): The HH-RLHF dataset contains human preference data for training language models to be helpful and harmless, as well as red teaming data to identify harmful model outputs. The preference data includes pairs of chosen and rejected responses, while the red teaming data includes transcripts of adversarial interactions with AI assistants, rated for harmfulness. We strictly follow prior works (Song et al., 2024; Yu et al., 2023) and used the code from this repository  [^2] https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/PRO for testing.
  • MATH (Hendrycks et al., 2021): The MATH dataset consists of 12,500 challenging competition-level math problems, each with a detailed step-by-step solution. It is designed to teach models to generate answer derivations and explanations, aiding in mathematical reasoning. Despite progress in improving accuracy, the dataset highlights the limitations of large Transformer models in solving complex math problems without new algorithmic advancements. Due to the high resource cost of using the full test set, we randomly sampled 400 problems from the test set for evaluation.
  • GSM8k (Chen et al., 2021): GSM8K (Grade School Math 8K) is a collection of 8.5K high-quality math word problems designed for grade school students. It supports the task of multi-step reasoning and question answering in basic math. The problems require 2 to 8 steps, focusing on elementary arithmetic operations (addition, subtraction, multiplication, and division). The solutions are provided in natural language, making it accessible for evaluation of language models’ internal reasoning. GSM8K has been widely used to test logic and mathematical capabilities in language models, especially for benchmarks like the LLM Leaderboard. Due to the high computational cost of using the entire test set, we randomly sampled 400 data points from the test set for our evaluation.
  • GAIA (Mialon et al., 2023): The GAIA dataset is a benchmark designed to evaluate next-generation LLMs with augmented capabilities like tooling and search access. It consists of over 450 complex questions with unambiguous answers, requiring various levels of autonomy and tooling. The dataset is divided into three levels, each increasing in difficulty, with a public dev set for validation and a private test set for evaluation. We used the entire test set for our evaluation.
  • HumanEval (Chen et al., 2021): The OpenAI HumanEval dataset contains 164 programming problems, each with a function signature, docstring, body, and unit tests. These problems are handwritten to ensure they were not included in the training sets of code generation models. The dataset is designed to evaluate the performance of models in Python code generation. We used the entire test set for evaluation.
  • AlpacaEval. (Dubois et al., 2024b): AlpacaEval consists of 805 instructions, including 252 from the self-instruct test set (Wang et al., 2022), 188 from the Open Assistant (OASST) test set, 129 from Anthropic’s helpful test set (Zhou et al., 2023), 80 from the Vicuna test set (Chiang et al., 2023), and 156 from the Koala test set (Vu et al., 2023).
  • LIMA. (Zhou et al., 2023): LIMA collects a training dataset of 1000 prompts and responses, curated to ensure stylistic consistency while accommodating diverse input types. It also includes an open-source test set of 300 prompts and a development set of 50. The data is primarily sourced from community-driven Q&A platforms like Stack Exchange, wikiHow, and the Pushshift Reddit Dataset (Baumgartner et al., 2020), along with manually curated examples. The inclusion of human-authored examples further increases dataset diversity. In our experiments, we use the LIMA test set to evaluate our models.
  • Vicuna. (Chiang et al., 2023): Vicuna organizes its 80 test instructions into eight distinct categories: Fermi problems, commonsense, roleplaying, coding/math/writing tasks, counterfactuals, knowledge, and general questions. This categorization aims to comprehensively assess different facets of chatbot performance. Prior work suggests that Vicuna’s instructions are generally of lower complexity and difficulty (Xu et al., 2023). We utilize the Vicuna test set to assess the performance of large language models across these varied categories of instructions.
  • Self-Instruct. (Wang et al., 2022): Self-Instruct contains 252 human-authored test instructions, each paired with a well-constructed output. This dataset is curated to simulate real-world use cases of instruction-following models, spanning various domains such as email composition, social media, productivity tools, and coding tasks. The instructions differ in format and complexity, featuring diverse task lengths and output types such as bullet points, tables, code snippets, and mathematical expressions. In our research, we utilized the Self-Instruct test set to rigorously evaluate our model’s ability to follow detailed instructions across multiple domains.
  • Wizardlm. (Xu et al., 2023): Wizardlm consists of a training set of 70k examples derived from 52k instructions initially provided by Alpaca. The test set includes 218 instructions sourced from various open-source projects and online communities, covering 29 distinct skills derived from real-world tasks. These skills range from Code Generation & Debugging to Reasoning, Mathematics, Writing, Complex Format Handling, and Mastery of Extensive Domains. In our study, we employed the Wizardlm test set to evaluate the model’s ability to adhere to detailed instructions comprehensively.
  • Koala. (Vu et al., 2023): Koala comprises 180 real-world user queries sourced from the web, spanning diverse topics and typically reflecting a conversational tone. These queries are especially relevant for evaluating models intended for chat-based applications. To ensure no overlap with training data, any query yielding a BLEU score above 20% compared to examples from our training set is excluded. Additionally, queries involving programming or non-English languages are omitted, as our evaluation team, composed of crowd-sourced raters, lacks the expertise to assess such content effectively. We exclusively used the Koala test set to gauge our model’s proficiency in handling authentic conversational queries.
A.8 Aggregation Process in GED Across Different Tasks
The GED implementation differs between response selection and model ranking tasks due to their objectives. In response selection, GED aggregates preference graphs from multiple evaluators for each question into a single DAG. From this DAG, a final ranking is derived, and the top-ranked answer is selected as the output. Notably, aggregation is performed only at the per-question level, without a rank aggregation across questions. In model ranking, GED involves two steps: first, aggregating the evaluators’ preference graphs into a DAG for each question to rank the models, and second, employing a rank ensemble method to aggregate these per-question rankings into a final overall model ranking across all questions. For details on the aggregation modeling, see Section 3.3, covering Response Selection, Model Ranking, and Model Alignment.
Appendix B Alternative Evaluator Configurations for Model Ranking

### v4

@@ 8 Data Selection for Model Fine-tuning
From Table 3, we observe that PGED consistently outperforms all baseline methods, demonstrating its effectiveness in selecting high-quality responses when multiple answers are available for a given instruction. When faced with multiple responses $y_{1},y_{2},\ldots,y_{n}$ for a given instruction $x$, the Random selection method can have a detrimental impact, especially when the quality of the responses is inconsistent. This effect is most evident with the Mistral-7B, where Random selection actually performs worse than the Origin, indicating that randomly chosen data points can introduce noise and degrade the model’s performance. Moreover, we find that simply selecting the longest response does not always lead to the best outcomes. While longer responses may provide more detailed answers, they are not necessarily better in terms of quality, particularly when both high-quality and low-quality answers exist for the same question. This is reflected in the results where the Longest method underperforms compared to both ContraSolver and PGED, emphasizing that response length alone is not always a reliable criterion. From Figure 3, we observe similar trends as in Table 3. PGED consistently outperforms all baselines across various datasets, demonstrating its effectiveness in selecting high-quality responses. Notably, in AlpacaEval and Self-Instruct, the Random baseline performs worse than the Origin model, highlighting that when response quality varies significantly, poor selection can degrade model performance. In contrast, PGED leverages preference graphs and denoising techniques to filter out low-quality responses, ensuring more robust and reliable performance, particularly in settings with inconsistent responses, as it removes evaluation noise and leads to more robust performance.
9 Case Study
(a) Raw Preference Graph 1
(b) Denoised Preference Graph 1
(c) Raw Preference Graph 2
(d) Denoised Preference Graph 2
Figure 4: Case studies showcasing the raw and denoised preference graphs.
In this section, we present two HumanEval case studies illustrating PGED on evaluator-aggregated graphs. Raw graphs are built from Llama3-8B, Mistral-7B, and Qwen2-7B judging ten Qwen2-72B responses (nodes 0–9). Figure 4 contrasts raw graphs (a, c) with denoised DAGs (b, d); dashed edges indicate removals. In Case 1, the raw graph contains cycles (e.g., $9\succ 4\succ 3\succ 9$, $9\succ 4\succ 3\succ 2\succ 9$). Removing a single edge $4\to 3$ breaks all cycles, yielding the DAG in Figure 4(b). In Case 2, cycles such as $7\succ 0\succ 6\succ 7$ and $7\succ 1\succ 0\succ 6\succ 7$ are eliminated by deleting $0\to 6$, producing the DAG in Figure 4(d). In both cases, alternative single-edge deletions fail to restore acyclicity without further removals. PGED thus identifies a minimal feedback arc set, preserving information while ensuring consistency.
10 Conclusion
In this paper, we presented PGED, a framework designed to address inconsistencies in pairwise preference evaluations by LLMs. By employing graph ensemble techniques and denoising, PGED reduces cyclic patterns and enhances the reliability of evaluation outcomes. Our theoretical analysis shows that PGED can recover the ground truth DAG under reasonable conditions, improving consistency in preference rankings. Extensive experiments across response selection, model ranking, and instruct tuning demonstrate the efficacy of our method. PGED consistently outperformed baseline methods in both single-evaluator and multi-evaluator settings, particularly in scenarios where combining small evaluators led to superior results over larger individual evaluators. Future work will explore extending PGED to broader evaluation frameworks and applying its principles to more complex decision-making tasks, including multi-agent systems and human-AI interaction.
Appendix A Implementation Details
A.1 Experimental Setup
All experimental procedures were conducted on a machine equipped with an AMD EPYC 7543 32-Core Processor, 512GB memory, 128 CPUs, and four 80GB NVIDIA A800 GPUs. The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Mistral-8×7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-7B-Instruct-v0.3, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct. We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al. 2024; Yu et al. 2023). Each experiment was repeated three times, and the average performance was reported as the final result. Our training script was adapted from the example provided in LlamaFactory  (Zheng et al. 2024b). The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and the model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler. For generating diverse responses from LLMs, we followed the configuration in (Yuan et al. 2024), setting $T=0.7$ and $p=0.9$. For tasks such as AlpacaEval (Dubois et al. 2024b), we used GPT-4-o unless stated otherwise.
A.2 Details of Evaluator Selection Across Different Tasks
Response Selection.
In the response selection task, we evaluated both single models and ensembles of evaluators using the PGEDmethod. For single model evaluation, we assessed the standalone performance of Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B on benchmarks such as HumanEval, AlpacaEval, MATH, GSM8k, and GAIA. This served two purposes: first, to establish the baseline performance of smaller models (Llama3-8B, Mistral-7B, Qwen2-7B) that are later combined using PGED, and second, to compare against Qwen2-72B, where we tested a baseline approach by randomly selecting one response from the 10 it generated for each question. For single evaluator evaluation, each model (Llama3-8B, Mistral-7B, Qwen2-7B, and Qwen2-72B) was also used as an evaluator to rank responses generated by Qwen2-72B. Notably, Qwen2-72B acted as a self-evaluator, selecting the best response from its own generated outputs. Finally, for multiple evaluator evaluation, our PGEDmethod combined the evaluations of Llama3-8B, Mistral-7B, and Qwen2-7B to rank responses.
Model Ranking.
For the model ranking task, we selected larger models as evaluators to ensure alignment with rankings produced by GPT-4. Specifically, we used Llama3-70B, Qwen2-72B, Qwen1.5-72B, and Mistral-8×7B as single evaluators. This choice was guided by two factors: performance considerations and practical feasibility. Larger models generally produce more reliable rankings, closely aligning with GPT-4, and the AlpacaEval benchmark, containing 805 tasks, makes the computational cost of using larger models acceptable. In the multiple evaluator setting, our PGEDmethod aggregated the evaluations from these four larger models to produce robust and consistent rankings. The combination of these high-capacity models ensures that our approach yields rankings that are both accurate and consistent across tasks.
Instruction Tuning.
In the instruction tuning task, the objective was to perform data selection for model training. For this task, we employed Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators. These models were selected because they balance computational efficiency and evaluation performance, making them suitable for iterative instruction tuning processes. The evaluators’ pairwise preferences were aggregated using PGEDto identify the most appropriate responses for training. This approach ensured that the selected data pairs reflected a consensus among the evaluators while keeping computational costs manageable.
A.3 Evaluation Settings: Single Model vs. Single Evaluator
The "single model" setting evaluates each model’s (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) outputs directly on benchmarks like HumanEval, AlpacaEval, MATH, GSM8k, and GAIA, without selection or modification. In contrast, the "single evaluator" setting uses these models to select the best response from ten candidates generated by Qwen2-72B, assessing their evaluation capability. The key difference is that the single model setting focuses on generation quality, while the single evaluator setting assesses evaluation ability.
A.4 Definition of Cycle Rate
A.5 Models Used for Ranking
We evaluate model ranking using 30 widely used models from the AlpacaEval dataset, covering diverse architectures and capabilities. Our model set $\mathcal{M}$ includes families such as Llama (Touvron et al. 2023), Vicuna (Zheng et al. 2023), GPT (Achiam et al. 2023), Claude (Anthropic 2024), Qwen (Yang et al. 2024), Mistral (Jiang et al. 2023), Yi (Young et al. 2024), and WizardLM (Xu et al. 2023), ensuring a broad assessment. Proprietary models include OpenAI’s GPT series (e.g., gpt-3.5-turbo-0301, gpt-4o-2024-05-13) and Anthropic’s Claude models (e.g., claude-2, claude-3-opus-20240229), serving as strong baselines. Open-source models include multiple versions of Llama (e.g., Meta-Llama-3-8B-Instruct, llama-2-70b-chat-hf) and Qwen (e.g., Qwen1.5-72B-Chat, Qwen2-72B). We also incorporate Mistral models (e.g., Mistral-7B-Instruct-v0.2, Mixtral-8x22B-Instruct-v0.1), reflecting advances in mixture-of-experts architectures. Additional selections include gemini-pro, tulu-2-dpo-70b, oasst-sft-llama-33b, dbrx-instruct, wizardlm-13b, and Yi-34B-Chat, contributing to ranking diversity. This diverse model set enables a comprehensive evaluation of ranking accuracy using AlpacaEval’s ground-truth rankings.
A.6 Dataset
  • UltraFeedback (Cui et al. 2023): UltraFeedback is a large-scale, fine-grained, diverse preference dataset, used for training powerful reward models and critic models. We collect about 64k prompts from diverse resources (including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN). We then use these prompts to query multiple LLMs (see Table for model lists) and generate 4 different responses for each prompt, resulting in a total of 256k samples.
  • HH-RLHF (Bai et al. 2022): The HH-RLHF dataset contains human preference data for training language models to be helpful and harmless, as well as red teaming data to identify harmful model outputs. The preference data includes pairs of chosen and rejected responses, while the red teaming data includes transcripts of adversarial interactions with AI assistants, rated for harmfulness. We strictly follow prior works (Song et al. 2024; Yu et al. 2023) for testing.
  • MATH (Hendrycks et al. 2021): The MATH dataset consists of 12,500 challenging competition-level math problems, each with a detailed step-by-step solution. It is designed to teach models to generate answer derivations and explanations, aiding in mathematical reasoning. Despite progress in improving accuracy, the dataset highlights the limitations of large Transformer models in solving complex math problems without new algorithmic advancements. Due to the high resource cost of using the full test set, we randomly sampled 400 problems from the test set for evaluation.
  • GSM8k (Chen et al. 2021): GSM8K (Grade School Math 8K) is a collection of 8.5K high-quality math word problems designed for grade school students. It supports the task of multi-step reasoning and question answering in basic math. The problems require 2 to 8 steps, focusing on elementary arithmetic operations (addition, subtraction, multiplication, and division). The solutions are provided in natural language, making it accessible for evaluation of language models’ internal reasoning. GSM8K has been widely used to test logic and mathematical capabilities in language models, especially for benchmarks like the LLM Leaderboard. Due to the high computational cost of using the entire test set, we randomly sampled 400 data points from the test set for our evaluation.
  • GAIA (Mialon et al. 2023): The GAIA dataset is a benchmark designed to evaluate next-generation LLMs with augmented capabilities like tooling and search access. It consists of over 450 complex questions with unambiguous answers, requiring various levels of autonomy and tooling. The dataset is divided into three levels, each increasing in difficulty, with a public dev set for validation and a private test set for evaluation. We used the entire test set for our evaluation.
  • HumanEval (Chen et al. 2021): The OpenAI HumanEval dataset contains 164 programming problems, each with a function signature, docstring, body, and unit tests. These problems are handwritten to ensure they were not included in the training sets of code generation models. The dataset is designed to evaluate the performance of models in Python code generation. We used the entire test set for evaluation.
  • AlpacaEval. (Dubois et al. 2024b): AlpacaEval consists of 805 instructions, including 252 from the self-instruct test set (Wang et al. 2022), 188 from the Open Assistant (OASST) test set, 129 from Anthropic’s helpful test set (Zhou et al. 2023), 80 from the Vicuna test set (Chiang et al. 2023), and 156 from the Koala test set (Vu et al. 2023).
  • LIMA. (Zhou et al. 2023): LIMA collects a training dataset of 1000 prompts and responses, curated to ensure stylistic consistency while accommodating diverse input types. It also includes an open-source test set of 300 prompts and a development set of 50. The data is primarily sourced from community-driven Q&A platforms like Stack Exchange, wikiHow, and the Pushshift Reddit Dataset (Baumgartner et al. 2020), along with manually curated examples. The inclusion of human-authored examples further increases dataset diversity. In our experiments, we use the LIMA test set to evaluate our models.
  • Vicuna. (Chiang et al. 2023): Vicuna organizes its 80 test instructions into eight distinct categories: Fermi problems, commonsense, roleplaying, coding/math/writing tasks, counterfactuals, knowledge, and general questions. This categorization aims to comprehensively assess different facets of chatbot performance. Prior work suggests that Vicuna’s instructions are generally of lower complexity and difficulty (Xu et al. 2023). We utilize the Vicuna test set to assess the performance of large language models across these varied categories of instructions.
  • Self-Instruct. (Wang et al. 2022): Self-Instruct contains 252 human-authored test instructions, each paired with a well-constructed output. This dataset is curated to simulate real-world use cases of instruction-following models, spanning various domains such as email composition, social media, productivity tools, and coding tasks. The instructions differ in format and complexity, featuring diverse task lengths and output types such as bullet points, tables, code snippets, and mathematical expressions. In our research, we utilized the Self-Instruct test set to rigorously evaluate our model’s ability to follow detailed instructions across multiple domains.
  • Wizardlm. (Xu et al. 2023): Wizardlm consists of a training set of 70k examples derived from 52k instructions initially provided by Alpaca. The test set includes 218 instructions sourced from various open-source projects and online communities, covering 29 distinct skills derived from real-world tasks. These skills range from Code Generation & Debugging to Reasoning, Mathematics, Writing, Complex Format Handling, and Mastery of Extensive Domains. In our study, we employed the Wizardlm test set to evaluate the model’s ability to adhere to detailed instructions comprehensively.
  • Koala. (Vu et al. 2023): Koala comprises 180 real-world user queries sourced from the web, spanning diverse topics and typically reflecting a conversational tone. These queries are especially relevant for evaluating models intended for chat-based applications. To ensure no overlap with training data, any query yielding a BLEU score above 20% compared to examples from our training set is excluded. Additionally, queries involving programming or non-English languages are omitted, as our evaluation team, composed of crowd-sourced raters, lacks the expertise to assess such content effectively. We exclusively used the Koala test set to gauge our model’s proficiency in handling authentic conversational queries.
A.7 Aggregation Process in PGEDAcross Different Tasks
The PGEDimplementation differs between response selection and model ranking tasks due to their objectives. In response selection, PGEDaggregates preference graphs from multiple evaluators for each question into a single DAG. From this DAG, a final ranking is derived, and the top-ranked answer is selected as the output. Notably, aggregation is performed only at the per-question level, without a rank aggregation across questions. In model ranking, PGEDinvolves two steps: first, aggregating the evaluators’ preference graphs into a DAG for each question to rank the models, and second, employing a rank ensemble method to aggregate these per-question rankings into a final overall model ranking across all questions. For details on the aggregation modeling, see Section 4, covering Response Selection, Model Ranking, and Model Alignment.
Appendix B Ablation study for Test Time scaling
We evaluate the impact of removing the ensembling step in PGED, referred to as the (w/o ensemble) variant. In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix Q). For simplicity of presentation, we use Weight Score to represent PGED(w/o ensemble) (Weight Score). As shown in Figure 5, all (w/o ensemble) methods consistently underperform compared to PGED. This performance gap arises because converting graphs to ranks before aggregation leads to information loss. In contrast, PGED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.
Figure 5: Comparison of PGED and (w/o ensemble) variants. PGED outperforms due to preserving more information by directly ensembling preference graphs, while rank aggregation in the (w/o ensemble) methods leads to performance loss.
Appendix C Limitation
Our framework improves consistency in preference evaluations, but it implicitly promotes a "majority view" by aggregating judgments from multiple LLM-based evaluators. This may inadvertently suppress minority or dissenting preferences, especially in cases where subjectivity, cultural variation, or ethical nuance plays a critical role. While the goal is to reduce noise, there is a risk that important human disagreement is treated as inconsistency to be "denoised." In socially sensitive domains, such as education, healthcare, or content moderation, this could limit the diversity of acceptable outputs or reinforce dominant norms embedded in current LLMs.
@@ Appendix C Limitation
Table 4: Performance comparison on nuanced quality metrics (%). PGEDoutperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful. and Validity metrics.
Method Fact. Rel. Coh. Info. Help. Valid. Avg.
@@
PGED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
Appendix D Evaluating PGED on More Metrics
@@
To assess PGEDon more metrics, we evaluated its impact on nuanced quality aspects of LLM outputs, including factuality, relevance, coherence, informativeness, helpfulness, and validity. Since LLM tasks often require subtle judgments, PGED’s adaptability is crucial. Following (Chen et al. 2023a), we used Llama3-70B to generate ten candidate responses per query and applied PGEDfor response selection, with Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators. Table 4 shows that PGEDconsistently outperformed individual evaluators and random selection across all metrics. It improved factuality by  5 % points over the best evaluator, while also enhancing relevance, coherence, informativeness, and helpfulness. These results highlight PGED’s ability to aggregate preferences effectively, capturing subtle qualities of generated content and improving overall reliability and utility in LLM assessments.
Appendix E Alternative Evaluator Configurations for Model Ranking